OpenStudy (anonymous):

Cost Function is c(x) = 2x+5 and the revenue function is R(x) = 2x - 2x^2 where x is the number of units (in thousands) produced and sold and R and C are measured in millions of dollars. Find: 1. Marginal Revenue 2. Marginal Cost 3. Break-even points 4. number of x for which marginal revenue equals marginal costs.

OpenStudy (anonymous):

1) Marginal revenue is the derivative of the revenue function, so take the derivative of R(x) and evaluate it at x 2) Marginal cost is the derivative of the cost function, so take the derivative and evaluate it at x 3) For break even set your cost formula equal to your revenue formula- because they should equal each other if you "break-even"

OpenStudy (anonymous):

What about this Solution....I am not sure people...can any one check, a) R(x)/x = 12 - 4x b) C(x)/x = 2 c) we want R(x) = C(x) 2x + 5 = 12x - 2x^2 12x - 2x^2-2x+5 2x^2-10x+5=0 4x-10 x = (5/2) +/- sqrt(60)/4 x = (5/2) +/- sqrt(15)/2 d) we want 12 - 2x > 2 + (5/x) multiply by x: 12x - 2x^2 > 2x + 5 2x^2 - 10x + 5 < 0 ie ((5/2) - sqrt(15)/2 ) < x < ((5/2) + sqrt(15)/2)

OpenStudy (anonymous):

Yes